# but Need a Good Lens

### v1

@@ Abstract
Previous research has primarily focused on the cognitive error detection capabilities of Large Language Models (LLMs), often prompting them to analyze mistakes in reasoning chains. However, few studies have examined the meta-cognitive abilities of LLMs (e.g., their self-awareness of step errors), which are crucial for their reliability. While studies on LLM self-evaluation present some measures, such as perplexity, which can reflect the answer correctness and be viewed as the lens of meta-cognition, they lack step-level analysis and adaptation. This paper studies the evaluation of LLM meta-cognition using the current lenses and how to improve these lenses. Specifically, we propose AutoMeco, an Automated Meta-cognition Evaluation framework for benchmarking the existing lenses. Furthermore, a training-free Markovian Intrinsic Reward Adjustment strategy, MIRA, is proposed to boost current meta-cognition lenses. Experimental results on three mathematical reasoning datasets and three LLMs show the reasonableness of AutoMeco by comparing it with Best-of-N verification. Moreover, the meta-cognition ability of LLMs can be better evaluated using MIRA. [^1] The code will be available upon publication.
@@
Ziyang Ma1, Qingyue Yuan2, Zhenglin Wang1, Deyu Zhou1† 1 School of Computer Science and Engineering, Key Laboratory of Computer Network and Information Integration, Ministry of Education, Southeast University, China 2 Department of Neurosurgery, Shanghai Tenth People’s Hospital, School of Clinical Medicine of Nanjing Medical University, China {mazy, d.zhou}@seu.edu.cn
22footnotetext: Corresponding Author.
@@ 3.2 AutoMated Meta-Cognition Evaluation
1 for each question $Q\in\mathcal{D}$ do
@@ 13                  $y_{i}^{\text{true}}\leftarrow 1$
Given an input question $Q$, the language model generates a response containing $N$ logically discrete reasoning steps $\{R_{i}\}_{i=1}^{N}$. Each step $R_{i}$ consists of consecutive tokens $[r_{1},r_{2},...,r_{{T_{i}}}]$ representing a coherent reasoning unit. We employ boundary detection based on transitional phrases (e.g., “Step 1:”, “Step 2:”, and “Answer:”) to segment the token sequence into interpretable steps.
@@ 8       $V(\mathcal{S}_{i})\leftarrow\max_{R_{i}}\mathcal{Q}(\mathcal{S}_{i},R_{i})$
The expected future reward $\mathcal{Q}(\mathcal{S}_{i},R_{i})$ integrates immediate confidence and discounted future value in Equation 9, with value function $V(\mathcal{S}_{i+1})=\max_{R_{i+1}}\mathcal{Q}(\mathcal{S}_{i+1},R_{i+1})$ and $\gamma\in(0,1]$ as the discount factor. From Equation 10, we recursively update Q-values from terminal state $\mathcal{S}_{N+1}$ with $V(\mathcal{S}_{N+1})=0$. This simplifies under a deterministic MDP to direct value assignment in Equation 11.
@@ 4.1 Experiments Setup
We conduct six meta-cognition lenses to predict the step correctness, which results in the intrinsic rewards of each step in the reasoning chains. These baselines are: (1) CoE-C Wang et al. (2025a); (2) CoE-R Wang et al. (2025a); (3) $\Delta$Entropy Yin et al. (2024); (4) Max Probability (Maxprob); (5) Perplexity (PPL) Huang et al. (2023); (6) Entropy Si et al. (2022). Among them, (1) and (2) require access to hidden states of LLMs, and (3)-(6) only require probability distributions. More details are in Appendix A.
@@
4.2 Statistical Feasibility
We prompt Qwen2.5-7B to reason on GSM8k, MATH500, and MinervaMATH. We conduct the self-evaluation measures to predict the step correctness, which results in the intrinsic rewards of each step in the reasoning chains. Subsequently, with the step rewards annotated by PRM, we calculate the correlation of the intrinsic and PRM rewards on the data split conditioned on PRM reward$\in(0,1]\cup[0.9,1)$. PRM denotes Qwen2.5-PRM-Math-7B in the following, if without reclarification. Besides, we show the distinguishability of step correctness by visualizing the kernel density estimation of Magnitude and Angle, two internal features proposed by Wang et al. (2025a). Appendix A presents details of the two features.
@@ 4.2 Statistical Feasibility
CoE-C 0.040 (0.174) 0.035 (0.074) 0.068 (0.0001) 0.046 (0.0001) 0.010 (0.710) 0.007 (0.685)
CoE-R 0.325 (1.7e-29) 0.225 (4.1e-30) 0.212 (1.4e-33) 0.145 (2.2e-34) 0.255 (3.7e-21) 0.175 (1.7e-21)
Maxprob 0.498 (1.3e-72) 0.354 (7.3e-72) 0.241 (2.1e-43) 0.164 (8.1e-44) 0.265 (9.6e-23) 0.182 (3.1e-23)
PPL 0.499 (6.1e-73) 0.355 (4.9e-72) 0.243 (4.3e-44) 0.165 (2.6e-44) 0.263 (2.2e-22) 0.180 (9.1e-23)
Entropy 0.521 (1.6e-80) 0.370 (3.1e-78) 0.246 (3.0e-45) 0.168 (1.1e-45) 0.270 (1.7e-23) 0.185 (6.8e-24)
$\Delta$Entropy -0.060 (0.041) -0.039 (0.052) 0.147 (7.5e-17) 0.097 (1.7e-16) 0.066 (0.017) 0.044 (0.016)
@@ 4.3 Comparison and Ablation Study
We evaluate the meta-cognition lenses in a more challenging and valuable setting with three language models. In this setting, we test the ability of these methods to distinguish the correct and incorrect steps, including wrong and unsure ones, instead of merely the wrong steps.
Table 2: AutoMeco and Best-of-N evaluation results of meta-cognition lenses with and without our proposed reward adjustment strategy, MIRA, across three mathematics reasoning datasets and three LLMs. Red and green denote whether MIRA improves the meta-cognition lenses. “Acc” denotes accuracy. “Majority” represents majority voting.
@@
Best-of-N Acc↑ AUROC↑ FPR95↓ AUPR↑ Best-of-N Acc↑ AUROC↑ FPR95↓ AUPR↑ Best-of-N Acc↑ AUROC↑ FPR95↓ AUPR↑
@@
Figure 3: Frequency of MIRA leading to improved, unchanged, and degraded performance for meta-cognition lenses on three LLMs and three datasets.
@@
Effectiveness of the Adjust Strategy
The proposed reward adjust strategy demonstrates robust generalization across model architectures and task difficulties. As detailed in Table 2, MIRA achieves performance improvements in 61.1% (BoN) and 68.5% (AUROC) of experimental configurations (N=54). Figure 3 illustrates the particular efficacy of MIRA in enhancing the meta-cognition lenses for Qwen2.5-7B. Although MIRA works on less than half of the settings for Llama-3-8B-Instruct, Table 4.3 presents that it enhances more meta-cognition lenses on the difficult task, MinervaMATH, than the simple one, GSM8k. Besides, MIRA also performs well on a wide range of meta-cognition lenses for Mistral-7B-Instruct. The gains brought by MIRA across different architectures prove its cross-model compatibility. These experimental results confirm our hypothesis that stepwise adjustment of intrinsic rewards can effectively compensate for inherent calibration weaknesses in the existing lenses of LLM meta-cognition. Although we primarily focus on open-source models, these results are still insightful for closed-source black-box LLMs due to the potential of probability-based meta-cognition lenses such as Maxprob and PPL.
@@
Practical validation on the MinervaMATH dataset reveals compelling advantages of meta-cognition lenses over conventional ensemble approaches. For instance, Qwen2.5-7B with adjusted CoE-R configuration outperforms majority voting baselines by 1.47% in BoN accuracy (12.13% vs 10.66%), while Llama-3-8B-Instruct achieves better performance through self-evaluation-based selection than majority voting (PPL: +0.36%, Entropy: +1.83%, $\Delta$entropy: +1.10%). This performance gap suggests that properly calibrated self-evaluation metrics enable more effective identification of high-quality reasoning paths than static aggregation methods. The findings align with our hypothesis that LLMs contain latent self-diagnostic capabilities that can be operationalized through appropriate metric design. Meanwhile, these results also indicate the potential of meta-cognition evaluation as a mechanism for autonomous LLM self-improvement through intrinsic process supervision.
@@ 4.4 PRM-as-a-Judge Analysis
To validate the reasonableness of utilizing PRM as an annotator, we compare AutoMeco with BoN by calculating the consistency of their benchmarking results of six meta-cognition lenses. The consistency metrics include top-K and last-K match rate, and consistency rate (CR) Wang et al. (2025b). Appendix B presents more details on these metrics.
Our analysis demonstrates that PRM-as-a-Judge achieves reasonable consistency with BoN evaluation, validating its utility as a complementary method for efficient method ranking. While perfect alignment is not observed, as expected given the inherent differences in evaluation mechanisms, the results highlight meaningful agreement trends.
As shown in Table 3, PRM-as-a-Judge exhibits alignment with BoN rankings for top-tier methods, particularly at K=3. BoN’s top methods include its top three in 66.67% settings on average, with Mistral achieving 100% consistency. Furthermore, the average CR of 48.15% supports the reasonableness of AutoMeco to benchmark the meta-cognition lenses. Moreover, it achieves the highest alignment on Llama-3 (CR=53.33%), suggesting model-specific optimization potential.
@@ 5 Conclusion
Our work focuses on introducing meta-cognition into language model evaluation by incorporating internal model states into self-assessment mechanisms. While our approach demonstrates promising results, two limitations warrant discussion:
@@ A.1 Chain-of-Embedding (CoE)
$\textrm{Mag}(\boldsymbol{H}_{i})$ $=\frac{1}{L}\sum_{\ell=0}^{L-1}\frac{||\boldsymbol{h}_{\ell+1}-\boldsymbol{h}_{\ell}||_{2}}{||\boldsymbol{h}_{L}-\boldsymbol{h}_{0}||_{2}}$
$\textrm{Ang}(\boldsymbol{H}_{i})$ $=\frac{1}{L}\sum_{\ell=0}^{L-1}\frac{\textrm{arccos}\left(\frac{\boldsymbol{h}_{\ell+1}^{\intercal}\ \cdot\ \boldsymbol{h}_{\ell}}{||\boldsymbol{h}_{\ell+1}||_{2}\ \cdot\ ||\boldsymbol{h}_{\ell}||_{2}}\right)}{\textrm{arccos}\left(\frac{\boldsymbol{h}_{L}^{\intercal}\ \cdot\ \boldsymbol{h}_{0}}{||\boldsymbol{h}_{L}||_{2}\ \cdot\ ||\boldsymbol{h}_{0}||_{2}}\right)}$
@@
$M(\boldsymbol{h}_{\ell},\boldsymbol{h}_{\ell+1})$ $=||\boldsymbol{h}_{\ell+1}-\boldsymbol{h}_{\ell}||_{2}$
$A(\boldsymbol{h}_{\ell},\boldsymbol{h}_{\ell+1})$ $=\textrm{arccos}\left(\frac{\boldsymbol{h}_{\ell+1}^{\intercal}\ \cdot\ \boldsymbol{h}_{\ell}}{||\boldsymbol{h}_{\ell+1}||_{2}\ \cdot\ ||\boldsymbol{h}_{\ell}||_{2}}\right)$
As shown in Equation 12 and 13, CoE-R and CoE-C capture the correctness by combining the magnitude and angle changes in the real and complex spaces, respectively Wang et al. (2025a).
@@ A.3 $\Delta$Entropy
We utilize the opposite number of $\Delta$Entropy as the step correctness score.
@@ Appendix B Consistency Metrics
[[[[ $\mathbb{I}[(\alpha_{m}-\alpha_{m^{\prime}})(\beta_{m}-\beta_{m^{\prime}})>0]-$

### v2

@@ Abstract
Previous research has primarily focused on the cognitive error detection capabilities of Large Language Models (LLMs), often prompting them to analyze mistakes in reasoning chains. However, few studies have examined the meta-cognitive abilities of LLMs (e.g., their self-awareness of step errors), which are crucial for their reliability. While studies on LLM self-evaluation present some measures, such as perplexity, which can reflect the answer correctness and be viewed as the lens of meta-cognition, they lack step-level analysis and adaptation. This paper studies the evaluation of LLM meta-cognition using the current lenses and how to improve these lenses. Specifically, we propose AutoMeco, an Automated Meta-cognition Evaluation framework for benchmarking the existing lenses. Furthermore, a training-free Markovian Intrinsic Reward Adjustment strategy, MIRA, is proposed to boost current meta-cognition lenses. Experimental results on three mathematical reasoning datasets and three LLMs show the reasonableness of AutoMeco by comparing it with Best-of-N verification. Moreover, the meta-cognition ability of LLMs can be better evaluated using MIRA. [^1] The code can be accessed via https://github.com/Yann-Ma/AutoMeco.
@@
Ziyang Ma1*, Qingyue Yuan2*, Zhenglin Wang1, Deyu Zhou1† 1 School of Computer Science and Engineering, Key Laboratory of Computer Network and Information Integration, Ministry of Education, Southeast University, China 2 Department of Neurosurgery, Shanghai Tenth People’s Hospital, School of Clinical Medicine of Nanjing Medical University, China {mazy, zhenglin, d.zhou}@seu.edu.cn  yuanqy007@stu.njmu.edu.cn
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding Author.
@@ 3.2 AutoMated Meta-Cognition Evaluation
1
21exfor each question $Q\in\mathcal{D}$ do
@@ 14       $y_{i}^{\text{true}}\leftarrow 1$
Given an input question $Q$, the language model generates a response containing $N$ logically discrete reasoning steps $\{R_{i}\}_{i=1}^{N}$. Each step $R_{i}$ consists of consecutive tokens $r_{1},r_{2},...,r_{{T_{i}}}]$ representing a coherent reasoning unit. We employ boundary detection based on transitional phrases (e.g., “Step 1:”, “Step 2:”, and “Answer:”) to segment the token sequence into interpretable steps.
@@ 9 $V(\mathcal{S}_{i})\leftarrow\max_{R_{i}}\mathcal{Q}(\mathcal{S}_{i},R_{i})$
The expected future reward $\mathcal{Q}(\mathcal{S}_{i},R_{i})$ integrates immediate confidence and discounted future value in Equation 9, with value function $V(\mathcal{S}_{i+1})=\max_{R_{i+1}}\mathcal{Q}(\mathcal{S}_{i+1},R_{i+1})$ and $\gamma\in(0,1]$ as the discount factor. In this equation, we recursively update Q-values from terminal state $\mathcal{S}_{N+1}$ with $V(\mathcal{S}_{N+1})=0$. This simplifies under a deterministic MDP to direct value assignment in Equation 10.
@@ 4.1 Experiments Setup
[…]
@@
We set the maximal number of new tokens for Qwen2.5-7B as 768 and those for the other two models as 2048 because of their extra instruction-tuning. In the AutoMeco evaluation, we set $\rm{temperature}=1.0$ and $\rm{num\_sequences}=1$ to evaluate the model under the condition of greedy generation. In the Best-of-N (BoN) evaluation, we set $\rm{temperature}=0.8$ and $\rm{N}=6$. We select the best response from the $N$ samples by choosing the one with the highest intrinsic reward averaged over steps. For Mistral-7B-Instruct, we use Mistral-7B-Instruct-v0.2 Jiang et al. (2023a). Appendix C shows the prompt templates for the three datasets. All the experiments are conducted on four 24G 3090 GPUs or one A100 80G GPU. For clear visualization, we use the kdeplot function from the seaborn libary Waskom (2021) with contour levels of 7, density thresholds of 0.15 for correct and 0.1 for incorrect samples, and $\textrm{bw\_adjust}=1.5$.
4.2 Statistical Feasibility
@@
CoE-C
0.040
0.035
0.068
0.046
0.010
0.007
CoE-R
0.325
(1.7e-29)
0.225
(4.1e-30)
0.212
(1.4e-33)
0.145
(2.2e-34)
0.255
(3.7e-21)
0.175
(1.7e-21)
Maxprob
0.498
(1.3e-72)
0.354
(7.3e-72)
0.241
(2.1e-43)
0.164
(8.1e-44)
0.265
(9.6e-23)
0.182
(3.1e-23)
PPL
0.499
(6.1e-73)
0.355
(4.9e-72)
0.243
(4.3e-44)
0.165
(2.6e-44)
0.263
(2.2e-22)
0.180
(9.1e-23)
Entropy
0.521
(1.6e-80)
0.370
(3.1e-78)
0.246
(3.0e-45)
0.168
(1.1e-45)
0.270
(1.7e-23)
0.185
(6.8e-24)
[…]
-0.060
-0.039
0.147
(7.5e-17)
0.097
(1.7e-16)
0.066
0.044
@@ 4.3 Comparison and Ablation Study
We evaluate the meta-cognition lenses in a more challenging and valuable setting with the three language models. In this setting, we test the ability of these methods to distinguish the correct and incorrect steps, including wrong and unsure ones, instead of merely the wrong steps.
Table 2: AutoMeco and Best-of-N evaluation results of meta-cognition lenses with and without our proposed reward adjustment strategy, MIRA, across three mathematics reasoning datasets and three LLMs. Green and red denote whether MIRA improves the meta-cognition lenses. “Acc” denotes accuracy. “Majority” represents majority voting. “PRM” denotes voting based on the step-averaged rewards calculated by the PRM.
@@
[…]
@@
Figure 3: Frequency of MIRA leading to improved, unchanged, and degraded performance for meta-cognition lenses on three LLMs (Qwen2.5-7B, Llama-3-8B-Instruct, Mistral-7B-Instruct) and three datasets (GSM8K, MATH500, MinervaMATH).
@@
Effectiveness and Efficiency of MIRA
The proposed reward adjust strategy demonstrates robust generalization across model architectures and task difficulties. As detailed in Table 2, MIRA achieves performance improvements in 61.1% (BoN) and 68.5% (AUROC) of experimental configurations (N=54). Figure 3 illustrates the particular efficacy of MIRA in enhancing the meta-cognition lenses for Qwen2.5-7B. Although MIRA works on less than half of the settings for Llama-3-8B-Instruct, Table 4.3 presents that it enhances more meta-cognition lenses on the difficult task, MinervaMATH, than the simple one, GSM8K. Besides, MIRA also performs well on a wide range of meta-cognition lenses for Mistral-7B-Instruct. The gains brought by MIRA across different architectures prove its cross-model compatibility. These experimental results confirm our hypothesis that stepwise adjustment of intrinsic rewards can effectively compensate for inherent calibration weaknesses in the existing lenses of LLM meta-cognition. Although we primarily focus on open-source models, these results are still insightful for closed-source black-box LLMs due to the potential of probability-based meta-cognition lenses such as Maxprob and PPL. Additionally, Table 3 demonstrates that MIRA adds only a negligible amount of latency per instance. The increase does not exceed 0.03 milliseconds, which validates the efficiency of MIRA.
Table 3: Latency measurements (unit: millisecond) for the meta-cognition lenses with (w/) and without (w/o) MIRA on Qwen2.5-7B (GSM8K subset). “Per Step” and “Per Instance” denote the latency averaged across all steps and all instances, respectively.
Methods Per Step Per Instance
w/o MIRA w/ MIRA
Maxprob 0.12 0.64 0.65
PPL 0.04 0.20 0.21
Entropy 512.94 2669.63 2669.64
[…]
CoE-R 0.51 2.64 2.66
CoE-C 0.64 3.31 3.34
@@
[…]
@@ 4.4 PRM-as-a-Judge Analysis
To validate the reasonableness of utilizing PRM as an annotator, we conduct the following analysis: (1) Analyze the consistency between the rewards of two different PRMs, Qwen2.5-Math-PRM-7B and Skywork-o1-Open-PRM-Qwen-2.5-7B He et al. (2024); (2) Compare AutoMeco with BoN by calculating the consistency of their benchmarking results of six meta-cognition lenses. The consistency metrics include top-K and last-K match rate, and consistency rate (CR) Wang et al. (2025b). Appendix B presents more details on these metrics.
Agreement between Different PRMs
Table 4 shows strong inter-annotator agreement between the two PRMs, confirming their reliability for evaluating step-level errors. Notably, the kappa score for annotations on Mistral-7B-Instruct’s reasoning traces reaches 0.7334. However, the PRMs exhibit lower agreement on step-level annotations than on instance-level ones. For example, the step-level Kappa for the reasoning steps of Qwen2.5-7B is 0.3274, compared to 0.5878 at the instance level. This discrepancy indicates that while the PRMs are effective judges overall, there is room to improve their fine-grained annotation accuracy.
Table 4: Inter-annotator agreement for step- and instance-level rewards between Qwen2.5-Math-PRM-7B and Skywork-o1-Open-PRM-Qwen-2.5-7B on the responses of the three LLMs to the three math datasets. “Spearman” and “Pearson” are the Spearman Spearman (1961) and Pearson’s correlation coefficients Pearson (1895). All the correlations in this table are statistically significant (p < 0.05). “Cohen” represents the inter-annotator agreement score, Cohen’s Kappa McHugh (2012). The reported Kappa score is maximized by finding the optimal binarization thresholds through a grid search over the range {0.1, 0.2, $\cdots$, 0.8, 0.9}.
[…]
Step Level
Qwen2.5-7B 0.6309 0.3552 0.3274
Llama-3-8B-Instruct 0.5850 0.3473 0.2657
Mistral-7B-Instruct 0.5758 0.6109 0.5349
Instance Level (Averaged on Steps)
Qwen2.5-7B 0.7630 0.6236 0.5878
Llama-3-8B-Instruct 0.7179 0.6617 0.6108
Mistral-7B-Instruct 0.7250 0.7981 0.7334
Consistency between AutoMeco and BoN
Our analysis demonstrates that PRM-as-a-Judge achieves reasonable consistency with BoN evaluation, validating its utility as a complementary method for efficient method ranking. While perfect alignment is not observed, the results highlight meaningful agreement trends. As shown in Table 5, AutoMeco exhibits alignment with BoN rankings for top-tier methods, particularly at K=3. BoN’s top methods include its top three in 66.67% settings on average, with Mistral-7B-Instruct achieving 100% consistency. Furthermore, the average CR of 48.15% supports the reliability of AutoMeco to benchmark the meta-cognition lenses. Moreover, it achieves the highest alignment on Llama-3-8B-Instruct (CR=53.33%), suggesting model-specific optimization potential.
@@ 5 Conclusion
Our work focuses on introducing meta-cognition into language model evaluation by incorporating internal model states into self-assessment mechanisms. While our approach demonstrates promising results, several limitations warrant discussion:
Model Accessibility Requirements
Our method incorporates internal model representations such as hidden states to enable fine-grained self-evaluation. This design, however, requires access to model internals and is mainly limited to open-source architectures. While logits- and probability-based self-evaluation methods are compatible with both open-source and mainstream closed-source LLMs with access to the last-layer states, more delicate approaches using hidden states cannot be directly applied to closed-source models such as GPT-4 Achiam et al. (2023). Exploring approximations or hybrid strategies may help bridge the gap between white-box and black-box interpretability.
@@
Acknowledgments
The authors would like to thank all of the anonymous reviewers for their valuable comments. This research is funded by the National Natural Science Foundation of China (Grant No.62176053). This work is supported by the Big Data Computing Center of Southeast University.
@@ A.1 Chain-of-Embedding (CoE)
[…]
@@
As shown in Equation 11 and 12, CoE-R and CoE-C capture the correctness by combining the magnitude and angle changes in the real and complex spaces, respectively Wang et al. (2025a).
$\textrm{CoE-R}(\boldsymbol{H}_{i})=\ \textrm{Mag}(\boldsymbol{H}_{i})-\textrm{Ang}(\boldsymbol{H}_{i})=\frac{1}{L}\sum_{\ell=0}^{L-1}\left(\frac{M(\boldsymbol{h}_{\ell},\boldsymbol{h}_{\ell+1})}{M(\boldsymbol{h}_{0},\boldsymbol{h}_{L})}-\frac{A(\boldsymbol{h}_{\ell},\boldsymbol{h}_{\ell+1})}{A(\boldsymbol{h}_{0},\boldsymbol{h}_{L})}\right)$ (11)
$\textrm{CoE-C}(\boldsymbol{H}_{i})$ $=\frac{1}{L}\left|\sum_{\ell=0}^{L-1}M(\boldsymbol{h}_{\ell},\boldsymbol{h}_{\ell+1})\cdot e^{i\cdot A(\boldsymbol{h}_{\ell},\boldsymbol{h}_{\ell+1})}\right|$
@@ A.2 Entropy
[…]
@@ Appendix B Consistency Metrics
$[$ $\mathbb{I}[(\alpha_{m}-\alpha_{m^{\prime}})(\beta_{m}-\beta_{m^{\prime}})>0]$-